. Three point charges are located in the x-y plane. q1 = -42.6 nC at (0.00, +3.60 cm), q2 = +45.4 nC at (-6.80 cm, 0.00) and q3 = -28.4 nC at (+8.20 cm, -9.40 cm). Determine the x component of the net electric field at the origin (0.00 m, 0.00m), due to the charges. Give your answer in the form "[+/-a.bc x 10^(x) N/C]i". Note that i is the unit vector in the x direction.
Answer #1


E1x = 0


E1y = +k*q1/y1^2 = +(9*10^9*42.6*10^-9)/(0.036^2) = 295833.33 N/C

E2x = +k*q2/x2^2 = +(9*10^9*45.4*10^-9)/(0.068^2) = -88365.052 N/C

E2y = 0

r3 = sqrt(x3^2+y3^2) = sqrt(0.082^2+0.094^2) = 0.125 m

E3x = k*q3*x3/r3^3

E3x = +(9*10^9*28.4*10^-9*0.082)/0.125^3= 10731.11 N/C

E3y = -k*q3*y3/r3^3


E3y = -(9*10^9*28.4*10^-9*0.094)/0.125^3 = -12301.52 N/C

Ex = E1x + E2x + E3x

Ex = -88365.052 + 10731.1 = -7.76*10^4 N/C

Ey = 295833.33 -12301.52 = 28.3*10^4 N/C

E = -7.76*10^4 i + 28.3*10^4 j